# Spike No More: Stabilizing the Pre-training of Large Language Models

### v2

@@ 2.2 Initialization
Because an initialization method has considerable influence on the training dynamics of neural methods, various initialization methods have been proposed for Transformers (Zhang et al., 2019b; Nguyen & Salazar, 2019; Zhang et al., 2019a; Wang et al., 2022). In this study, we investigate the training stability of Transformers in two kinds of initialization methods; we call them plain initialization and scaled initialization. In the plain initialization, we initialize all parameters with a normal distribution $\mathcal{N}(0,\sigma)$. In the scaled initialization, which is the widely used method in the pre-training of LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023), we scale $W_{2}$ and $W_{O}$ to small values. Concretely, we follow Shoeybi et al. (2020); Le Scao et al. (2022) and initialize $W_{2}$ and $W_{O}$ with a normal distribution $\mathcal{N}(0,\sigma/\sqrt{2N})$ where $N$ is the number of layers [^2] Biderman et al. (2023) also scaled $W_{2}$ and $W_{O}$ to small values in the initialization, but they used the strategy introduced by Wang & Komatsuzaki (2021) instead of $\mathcal{N}(0,\sigma/\sqrt{2N})$. We discuss the effect of this initialization method in Appendix C.. Following Le Scao et al. (2022), we use $\sqrt{2/5d}$ for $\sigma$, that is initially proposed by Nguyen & Salazar (2019).
3 Analyses on Exploding Gradients
This section discusses two causes of the exploding gradients and the ways to prevent them. We indicate that the explosion due to the shortcut occurs when we use the plain initialization described in Section 2.2. Moreover, we indicate that the scaled initialization mitigates the explosion due to the shortcut but causes the explosion due to LNs. To prevent both explosions, we describe a requirement and introduce a modification in the scaled initialization.
3.1 Explosion due to Shortcut
In this section, we indicate that the exploding gradients due to the shortcut occur when we use the plain initialization. Moreover, we indicate that the scaled initialization prevents the exploding gradients due to the shortcut.
For the discussion, we analyze the norms of each layer output calculated by Equations (1) and (2). We assume that the vector $x$ at each layer follows the normal distribution, and then we can assume that the mean of the vector $x$ is 00. In this case, we have to consider only the standard deviation of the vector $x$ to estimate its norm because the standard deviation is in proportion to the norm. In detail, let $\mathrm{std}(x)$ be the standard deviation of $x$, $||x||_{2}=\sqrt{d}\ \mathrm{std}(x)$. In addition, let $\mathrm{var}(x)$ represent the variance of $x$, we obtain the variance of the sum of two vectors that follow the normal distribution with the following equation:
$\mathrm{var}(x+y)=\mathrm{var}(x)+\mathrm{var}(y).$ (6)
For both initialization methods, we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma=\sqrt{2/5d}\ll 1$, and thus, the standard deviations of embeddings are much smaller than $1$. Then, we investigate the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ in Equations (3) and (4). If these standard deviations are also much smaller than $1$, the standard deviations of each layer output by Equations (1) and (2) remain on a small scale. In contrast, for the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$, if either is close to or larger than $1$, the standard deviation of the shortcut rapidly grows as the layer gets deeper due to Equation (6). This leads to the exponential growth of the norms of each layer calculated by Equations (1) and (2), and thus, causes exploding gradients during the back-propagation.
Let consider $\mathrm{var}(\mathrm{FFN}(x))$. To simplify $\mathrm{FFN}(x)$ in Equation (3), we assume that $\mathcal{F}$ is the identity function. Then, we obtain $\mathrm{var}(\mathrm{FFN}(x))$ at the beginning of the training with the following equation:
$\mathrm{var}(\mathrm{FFN}(x))=d_{ffn}\ d\ \mathrm{var}(x)\ \mathrm{var}(W_{1})\ \mathrm{var}(W_{2}).$ (7)
Here, $\mathrm{var}(x)=1$ because we apply LN before the calculation of $\mathrm{FFN}$ as in Equation (1). We set $d_{ffn}=4d$ as the widely used configuration. For $\mathrm{var}(W_{1})$ and $\mathrm{var}(W_{2})$, we assign actual values described in Section 2.2. For example, in the plain initialization,
$\mathrm{var}(\mathrm{FFN}(x))$ $=4d\ d\ \frac{2}{5d}\frac{2}{5d}$
$=\frac{16}{25}.$ (8)
Thus, the standard deviation is $\sqrt{16/25}=4/5\simeq 1$. Because at least the standard deviation of $\mathrm{FFN}(x)$ is close to $1$, the plain initialization exponentially increases the standard deviation of the shortcut during forward computation, and then causes exploding gradients during the back-propagation.
In contrast, the scaled initialization alleviates the explosion by scaling $W_{2}$ to smaller values, $1/\sqrt{2N}$. Introducing the scaling, we obtain the following equation:
$\mathrm{var}(\mathrm{FFN}(x))$ $=4d\ d\ \frac{2}{5d}\frac{2}{5d}\frac{1}{2N}$
$=\frac{8}{25N}.$ (9)
Thus, the standard deviation is $\sqrt{8/25N}$. Recall that $N$ represents the number of layers as defined in Section 2.2, $\sqrt{8/25N}$ is much smaller than 1 for typical LLMs since they usually use $N>10$.
We indicated that the scaled initialization mitigates the increase of the shortcut by $\mathrm{FFN}(x)$. We prove that the scaled initialization also mitigates the increase by $\mathrm{Attention}(x)$. Let us consider the standard deviation of Equation (4) in the scaled initialization. Before computing, we consider the standard deviation of each self-attention head computed by Equation (5). To simplify the equation, let $\frac{xW_{Qi}(XW_{Ki})^{\mathrm{T}}}{\sqrt{d_{head}}}$ be $Z$. Then, we can transform the right term of Equation (5) to $\mathrm{softmax}(Z)XW_{Vi}$. Because $X$ is the pack of layer-normalized vectors, the variance of $\mathrm{softmax}(Z)X$ reaches the maximum value $1$ when the result of $\mathrm{softmax}(Z)$ is a $1$-hot vector [^3] On the other hand, the variance reaches the minimum value when the result of the $\mathrm{softmax}$ is a uniform distribution.. Thus, we obtain the variance of Equation (5) as follows:
$\mathrm{var}(\mathrm{head}_{i}(x))$ $=\mathrm{var}(\mathrm{softmax}(Z)X)\ d\ \mathrm{var}(W_{Vi})$
$=\mathrm{var}(\mathrm{softmax}(Z)X)\ d\ \frac{2}{5d}$
$<\frac{2}{5}.$ (10)
Therefore, since we also scale $W_{O}$ to $1/\sqrt{2N}$, the variance of Equation (4) is as follows:
$\mathrm{var}(\mathrm{Attention}(x))$ $=\mathrm{var}(\mathrm{head}(x))\ d\ \mathrm{var}(W_{O})$
$=\mathrm{var}(\mathrm{head}(x))\ d\ \frac{2}{5d}\frac{1}{2N}$
$=\mathrm{var}(\mathrm{head}(x))\ \frac{1}{5N}$
$<\frac{2}{25N}.$ (11)
Based on these calculations, the standard deviation of $\mathrm{Attention}(x)$ is also much smaller than $1$ when we use the scaled initialization. These theoretical analyses prove that the scaled initialization prevents the rapid growth of the shortcut, and thus, it prevents the exploding gradients due to the shortcut.
Figure 2: Gradient norms of Transformers at the beginning of the LLM pre-training. To focus on only the explosion due to the shortcut, we prevent the explosion due to LNs by multiplying embeddings by $\sqrt{d}$ as described in Section 3.2.
Figure 2 shows the gradient norms of each layer at the beginning of the LLM pre-training. This figure shows that the gradient norms in the plain initialization rapidly grow as the layer gets shallower. This means that the plain initialization faces the exploding gradients. In contrast, the gradient norms in the scaled initialization remain at almost constant values. Therefore, the scaled initialization, which is widely used in LLM pre-training, is effective in preventing the exploding gradients.
However, in this figure, we prevent the explosion due to LNs to focus on only the explosion due to the shortcut. If we remove the modification, the gradient norms in the scaled initialization also explode. We describe the explosion due to LNs and the modification to address it in the next section.
3.2 Explosion due to LNs
Xiong et al. (2020) proved that the gradient norm of the LN depends on the norm of an input vector as follows:
$\left\lVert\frac{\delta\mathrm{LN}(x)}{\delta x}\right\rVert_{2}=\mathcal{O}\left(\frac{\sqrt{d}}{||x||_{2}}\right).$ (12)
As described in Section 3.1, we assume $||x||_{2}=\sqrt{d}\ \mathrm{std}(x)$. Thus, we can modify the above equation as follows:
$\left\lVert\frac{\delta\mathrm{LN}(x)}{\delta x}\right\rVert_{2}=\mathcal{O}\left(\frac{1}{\mathrm{std}(x)}\right).$ (13)
Therefore, the LN increases its gradient norm drastically when $\mathrm{std}(x)\ll 1$. We investigate whether the standard deviations of vectors before each LN are smaller than $1$ to find the point where the gradient increases.
The input vectors of LNs are an embedding at the first layer, and each layer output that is the sum of the shortcut and $\mathrm{FFN}(x)$ or $\mathrm{Attention}(x)$ as in Equations (1) and (2). As described, the standard deviations of embeddings are much smaller than $1$ because we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma\ll 1$. In addition, since the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ are much smaller than $1$ in the scaled initialization as described in Section 3.1, LNs certainly increase gradients due to Equation (13) especially in shallow layers [^4] The each layer output grows as the layer gets deeper, because we add $\mathrm{FFN}(x)$ and $\mathrm{Attention}(x)$ the number of layers times. For example, we can remove $N$ in denominators of Equations (9) and (11) at the final layer. Therefore, the standard deviations of each layer output at deep layers are close to $1$ in comparison with ones at shallow layers..
To solve this issue, we need to make the standard deviations of vectors immediately before LNs close to $1$. We have two choices to achieve this requirement. One is increasing the standard deviation of the shortcut, and the other is increasing the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention(x)}$. However, as described in Section 3.1, if we make the standard deviations of $\mathrm{FFN}(x)$ and $\mathrm{Attention(x)}$ close to $1$, the explosion due to the shortcut occurs because the norm of the shortcut rapidly grows. Thus, we make the standard deviation of the shortcut close to $1$.
@@ 3.2 Explosion due to LNs
To make the standard deviation of the shortcut close to $1$, we make the standard deviation of each embedding close to $1$. To achieve this, we introduce two kinds of modification: “Scaled Embed” and “Embed LN”. The Scaled Embed is scaling embeddings with an appropriate value [^5] In this study, we use word tying (Press & Wolf, 2017; Inan et al., 2017). In this configuration, we have to initialize embeddings with a small value to use them in the output layer. Otherwise, we can initialize embeddings with the normal distribution $\mathcal{N}(0,\sigma)$ where $\sigma=1$ to solve the explosion due to LNs.. For example, we multiply embeddings by $\sqrt{d}$, which is introduced in the original Transformer paper (Vaswani et al., 2017) [^6] Although the original Transformer paper introduced this operation, recent implementations ignore this., and then the standard deviations of embeddings become $\sqrt{2/5}$. The Embed LN is applying the LN to embeddings. In fact, Le Scao et al. (2022) reported that the Embed LN strategy prevents loss spikes empirically.
Figure 3 shows the gradient norms of each layer when we use the scaled initialization. This figure indicates that the gradient exponentially increases especially in shallow layers in Vanilla that uses the scaled initialization but does not use any modification to prevent the explosion due to LNs. In contrast, our modifications prevent the exploding gradients because they resolve the explosion due to LNs. Therefore, we require the combination of the scaled initialization and modifications described in this section to prevent both explosions due to the shortcut and LNs.
Figure 3: Gradient norms of Transformers at the beginning of the LLM pre-training. To focus on only the explosion due to LNs, we use the scaled initialization method described in Section 2.2.
@@ 4 Main Experiments
We verify the empirical effectiveness of our theoretical analyses. In detail, we demonstrate that the method that prevents the exploding gradients also prevents loss and gradient spikes. In particular, we investigate the effectiveness of modifications to prevent the explosion due to LNs because the empirical effectiveness of the scaled initialization has already been revealed in previous studies (Shoeybi et al., 2020; Le Scao et al., 2022).
@@
4.1 Datasets
4.2 Model Configurations
As described in Section 2, we used the Pre-LN Transformer architecture. In addition, we use the scaled initialization because the method is widely used in LLM pre-trainings (Shoeybi et al., 2020; Le Scao et al., 2022). We set the number of layers $N=24$, and varied $d$ to adjust the total number of parameters to 350 million (350M) and 1.7 billion (1.7B). We set the learning rate ($\mathrm{lr}$) 5.0 ×$\times 10^{-4}$. Section 5.1 shows experiments in varying the learning rate. Appendix A describes more details on the experimental configuration.
We compared the following methods. We put $\checkmark$ before the method name if the method satisfies the requirements to prevent the explosion due to LNs.
@@ 4.2 Model Configurations
This configuration uses only the scaled initialization. Thus, this configuration prevents the explosion due to the shortcut but faces the explosion due to LNs.
@@
$\mathrm{Embed}=\gamma\mathrm{Embed}+(1-\gamma)\mathrm{Detach}(\mathrm{Embed}),$ (14)
where $\gamma$ is a hyper-parameter and $\mathrm{Detach}$ detaches an input from the computational graph. We assign $0.1$ to $\gamma$ as in Zeng et al. (2023). Zeng et al. (2023) indicated that this method empirically prevents the loss spike. However, this method does not satisfy the requirements to prevent the explosion due to LNs, and thus, we show that this method does not completely solve the loss spike.
@@
Dettmers et al. (2022) and Le Scao et al. (2022) reported that applying the LN to the embedding layer stabilizes their LLM pre-training. As described in Section 3.2, this method satisfies the requirements to prevent the explosion due to LNs in internal layers.
@@
This method multiplies embeddings by $\sqrt{d}$. As described in Section 3.2, this method satisfies the requirements to prevent the explosion due to LNs.
@@ 4.3 Results
Figure 4 shows the loss values of each method in validation data. Figure 5 shows the gradient norms of each method. These figures indicate that Vanilla and Embed Detach faced loss and gradient spikes. In contrast, Embed LN and Scaled Embed did not face spikes. These results correspond to our theoretical analyses described in Section 3. Thus, we must prevent the exploding gradients to avoid spikes in LLM pre-training.
In comparison between 350M and 1.7B parameters, spikes occurred more frequent in 1.7B parameters. Because we initialize embeddings with $\mathcal{N}(0,\sigma)$ where $\sigma=\sqrt{2/5d}$, the standard deviations of embeddings become small as $d$ gets larger. This means that the standard deviation of Equations (1) and (2) also become small as $d$ gets larger. As described in Section 3.2, the smaller the standard deviation the LN input is, the larger its gradient is. Therefore, if we increase $d$ without any technique to prevent the explosion due to LNs, a model becomes more unstable. This result corresponds to the previous study reports (Le Scao et al., 2022; Chowdhery et al., 2022; Zeng et al., 2023) that their model became more unstable as they increased the number of parameters.
Table 1: Perplexities of each method with 350M and 1.7B parameters.
@@
5 Difference from Other Configurations
@@ 5.1 Varying Learning Rate
Figure 6 shows loss values of each configuration in validation data. As shown in this figure, the larger the learning rate we used, the more frequent the spikes occurred in Vanilla. In particular, in $\mathrm{lr}=1.0\times 10^{-3}$, the training of Vanilla with 1.7B parameters failed because its gradient exploded, as shown in Figure 7. In contrast, Scaled Embed stabilized the training, and thus, its loss values consistently decreased.
Table 2 shows the perplexities of each configuration in evaluation data. This table indicates that Vanilla with 350M parameters achieved better performance in $\mathrm{lr}=1.0\times 10^{-4}$ that is the situation where its training did not face any spike. This result corresponds to the report of Le Scao et al. (2022). Thus, we suppose that they conducted the comparison with a too-small learning rate to stabilize Vanilla. In contrast, the stable methods are more effective in training with a large learning rate as shown in Figure 6 and Table 2. Therefore, if Le Scao et al. (2022) used a relatively large learning rate in their experiments, their stable method achieved better performance.
(a) 350M parameters with $\mathrm{lr}=1.0\times 10^{-3}$.
(b) 350M parameters with $\mathrm{lr}=5.0\times 10^{-4}$.
(c) 350M parameters with $\mathrm{lr}=1.0\times 10^{-4}$.
(d) 1.7B parameters with $\mathrm{lr}=1.0\times 10^{-3}$.
(e) 1.7B parameters with $\mathrm{lr}=5.0\times 10^{-4}$.
(f) 1.7B parameters with $\mathrm{lr}=1.0\times 10^{-4}$.
@@
Figure 7: Gradient norms of each method with $\mathrm{lr}=1.0\times 10^{-3}$.
@@
Model $\mathrm{lr}=1.0\times 10^{-3}$ $\mathrm{lr}=5.0\times 10^{-4}$ $\mathrm{lr}=1.0\times 10^{-4}$ $\mathrm{lr}=1.0\times 10^{-3}$ $\mathrm{lr}=5.0\times 10^{-4}$ $\mathrm{lr}=1.0\times 10^{-4}$

### v3

@@ 4 Conditions to Avoid Spikes
Based on the discussions in Section 3, we have to pay attention to values of $\sigma_{1}$, $\sigma_{2}$, $\sigma_{O}$, and the standard deviation of the shortcut to stabilize the pre-training of LLMs. To make $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{O}$ small, we have to initialize the corresponding parameters with a small value. Let us consider the actual settings in detail. The widely used initialization method for LLMs (Shoeybi et al., 2020; Le Scao et al., 2022; Biderman et al., 2023), initializes all parameters with a normal distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{\frac{2}{5d}}$ (Nguyen & Salazar, 2019), and then scales $W_{2}$ and $W_{O}$ to small values based on the number of layers: $\sqrt{\frac{1}{2N}}$ where $N$ is the number of layers [^5] Biderman et al. (2023) also scaled $W_{2}$ and $W_{O}$ to small values in the initialization, but they used the strategy introduced by Wang & Komatsuzaki (2021) instead of scaling with $\sqrt{\frac{1}{2N}}$. However, its property is the same essentially because they initialize $W_{2}$ and $W_{O}$ with $\sigma=\frac{2}{N\sqrt{d}}$ which becomes small based on the number of layers.. In this situation, $\sigma_{1}$, $\sigma_{2}$, and $\sigma_{O}$ are sufficiently small values.
@@
However, in this situation, the standard deviation of the shortcut is also too small. For example, at shallow layers, the standard deviation is close to $\sqrt{\frac{2}{5d}}$ because the embedding matrix is also initialized by $\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{\frac{2}{5d}}$. Therefore, to increase the standard deviation of the shortcut, we make the standard deviation of each embedding close to 1111666Based on Equations (15) and (20), the upper bound becomes small as the standard deviation of the shortcut increases. However, a too large value degrades the performance empirically as described in Appendix E.. To achieve this, we introduce two kinds of modification: “Scaled Embed” and “Embed LN” [^7] We can satisfy the condition by initializing embeddings with the normal distribution $\mathcal{N}(0,\sigma^{2})$ where $\sigma=1$, but we do not adopt this strategy in this study because we use the same initialization method in our experiments.. The Scaled Embed scales embeddings with an appropriate value. For example, we multiply embeddings by $\sqrt{d}$, which was used in the original Transformer paper (Vaswani et al., 2017) [^8] Although the original Transformer paper introduced this operation, recent implementations ignore this., and then the standard deviations of embeddings become $\sqrt{\frac{2}{5}}$. The Embed LN applies the LN to embeddings. In fact, Le Scao et al. (2022) reported that the Embed LN strategy prevents loss spikes empirically. These two methods are presented as verification examples rather than proposed methods, and alternative approaches could be employed if the conditions are met.
Figure 2: The actual upper bound described in Equation (15) for each Transformer layer at the beginning of the LLM pre-training. Because it is difficult to estimate the strict values for $\sigma_{x}$ at all layers, we obtain the empirical values by using some inputs, and assign them to Equation (15).
To demonstrate the actual values of the upper bound described in Equation (15), we take the model with 1.7 billion parameters as an example. In addition to the widely used initialization for LLMs (Vanilla) and the above two modifications: Scaled Embed and Embed LN, we compare Xavier Init, which initializes all parameters with the Xavier initialization (Glorot & Bengio, 2010), as the situation where we do not scale $W_{2}$ and $W_{O}$ based on the number of layers. Figure 2 shows the values of Equation (15) for each layer at the beginning of the pre-training. This figure indicates that the methods without suppressing the upper bound, i.e., Xavier Init and Vanilla, rapidly increase the values especially in shallow layers. In contrast, Scaled Embed and Embed LN keep small values. In summary, to make the upper bound of the gradient norms small for the stabilization of the LLM pre-training, we have to satisfy two conditions: (1) small sub-layers; initializing the parameters of sub-layers with a small value and (2) large shortcut; making the standard deviation of each embedding close to $1$.
Table 1: Relations between each method in experiments and two conditions to control the upper bound of gradient norms. We conduct experiments for Xavier Init and Xavier Init + Scaled Embed in Appendix B.
Method Small sub-layers Large shortcut
Xavier Init - -
Vanilla $\checkmark$ -
Embed Detach $\checkmark$ -
Embed LN $\checkmark$ $\checkmark$
Scaled Embed $\checkmark$ $\checkmark$
Xavier Init + Scaled Embed - $\checkmark$
@@ 5 Main Experiments
We verify the empirical effectiveness of our theoretical analyses. In detail, we demonstrate that controlling the upper bound of the gradient norms also prevents loss and gradient spikes. To assess efficacy in the real situation, we focus on the methods initialized with the widely used method (Shoeybi et al., 2020; Le Scao et al., 2022) in main experiments [^9] The Xavier initialization, which does not satisfy the small sub-layers condition as shown in Table 1, is not widely used for LLMs in recent years. Appendix B shows that the performance of Xavier initialization is worse and fails to avoid spikes..
5.1 Datasets
5.2 Model Configurations
As described in Section 2, we used the Pre-LN Transformer architecture. We set the number of layers $N=24$, and varied $d$ to adjust the total number of parameters to 350 million (350M) and 1.7 billion (1.7B). We set the learning rate ($\mathrm{lr}$) $5.0\times 10^{-4}$. Section 6.1 shows experiments in varying the learning rate. Appendix A describes more details on the experimental configuration.
@@ 5.2 Model Configurations
We compared the following methods. We put $\checkmark$ before the method name if the method satisfies both conditions to suppress the upper bound. Moreover, Table 1 summarizes whether each method satisfies each condition.
@@
This is the standard configuration for the LLM pre-training. Since this configuration does not suppress the upper bound of the gradient norms, the loss spike is likely to occur.
@@
$\mathrm{Embed}\leftarrow\gamma\mathrm{Embed}+(1-\gamma)\mathrm{Detach}(\mathrm{Embed}),$ (21)
where $\gamma$ is a hyper-parameter and $\mathrm{Detach}$ detaches an input from the computational graph. We assign $0.1$ to $\gamma$ as in Zeng et al. (2023). Zeng et al. (2023) indicated that this method empirically prevents the loss spike. However, this method does not satisfy the condition on large shortcut, and thus, we show that this method does not completely solve the loss spike.
@@
Dettmers et al. (2022) and Le Scao et al. (2022) reported that applying the LN to the embedding layer stabilizes their LLM pre-training. As described in Section 4, this method satisfies the conditions to control the upper bound of the gradient norms.
@@
This method multiplies embeddings by $\sqrt{d}$. As described in Section 4, this method satisfies the requirements to control the upper bound of the gradient norms.
@@ 5.3 Results
Figure 3 shows the loss values of each method in validation data. Figure 4 shows the gradient norms of each method. These figures indicate that Vanilla and Embed Detach faced loss and gradient spikes. In contrast, Embed LN and Scaled Embed did not face spikes. These results correspond to our theoretical analyses described in Sections 3 and 4. Thus, only methods that make the upper bound of the gradient norm small have successfully avoided spikes in LLM pre-training.
In comparison between 350M and 1.7B parameters, spikes occurred more frequently in 1.7B parameters. Because we initialize embeddings with $\mathcal{N}(0,\sigma^{2})$ where $\sigma=\sqrt{\frac{2}{5d}}$, the standard deviations of embeddings become small as $d$ gets larger in Vanilla and Embed Detach. This means that the upper bounds described by Equations (15) and (20) become large as $d$ gets larger because $\sigma_{x}$ and $\sigma_{x}^{\prime}$ are nearly equal to the standard deviation of an input embedding in shallow layers. Therefore, if we increase $d$ without any technique to control the upper bound of the gradient norms, a model becomes more unstable. This result corresponds to the previous study reports (Le Scao et al., 2022; Chowdhery et al., 2022; Zeng et al., 2023) that their model became more unstable as they increased the number of parameters.
Table 2: Perplexities of each method.
@@
6 Discussions on Other Configurations
@@ 6.1 Varying Learning Rate
(a) 350M with $\mathrm{lr}=1.0\times 10^{-3}$.
(b) 350M with $\mathrm{lr}=5.0\times 10^{-4}$.
(c) 350M with $\mathrm{lr}=1.0\times 10^{-4}$.
(d) 1.7B with $\mathrm{lr}=1.0\times 10^{-3}$.
(e) 1.7B with $\mathrm{lr}=5.0\times 10^{-4}$.
(f) 1.7B with $\mathrm{lr}=1.0\times 10^{-4}$.
@@
Figure 5 shows loss values of each configuration in validation data. As shown in this figure, the larger the learning rate we used, the more frequent the spikes occurred in Vanilla. In particular, in $\mathrm{lr}=1.0\times 10^{-3}$, the training of Vanilla with 1.7B parameters failed because its gradient exploded. In contrast, Scaled Embed stabilized the training, and thus, its loss values consistently decreased.
Table 3 shows the perplexities of each configuration in evaluation data. This table indicates that Vanilla with 350M parameters achieved better performance in $\mathrm{lr}=1.0\times 10^{-4}$ that is the situation where its training did not face any spike. This result corresponds to the report of Le Scao et al. (2022). Thus, we suppose that they conducted the comparison with a too-small learning rate to stabilize Vanilla. In contrast, the stable methods are more effective in training with a large learning rate, as shown in Figure 5 and Table 3. Therefore, if Le Scao et al. (2022) used a relatively large learning rate in their experiments, their stable method could achieve better performance.
@@
Model lr $\mathrm{lr}\ 1.0\times 10^{-3}$ lr $\mathrm{lr}\ 5.0\times 10^{-4}$ lr $\mathrm{lr}\ 1.0\times 10^{-4}$ lr $\mathrm{lr}\ 1.0\times 10^{-3}$ lr $\mathrm{lr}\ 5.0\times 10^{-4}$ lr $\mathrm{lr}\ 1.0\times 10^{-4}$